OpenStudy (anonymous):

Hi can you integrate (-2x^2-x) / (2x^3+x^2+1) with respect to x?

OpenStudy (anonymous):

you may need to decompose using the partial fractions technique

OpenStudy (anonymous):

ohh i thought u are going to show it up.

OpenStudy (anonymous):

but thanks for the tips...

OpenStudy (anonymous):

Hai UnkleRhaukus? Are you writing an answer?

OpenStudy (anonymous):

I will give the best answer to him. If not..

OpenStudy (unklerhaukus):

can you do some thing tricky and make the integrand du/u ?

OpenStudy (unklerhaukus):

hmm , it dosen't quite work

OpenStudy (anonymous):

is it possible to integrate ?

OpenStudy (anonymous):

6x^2+2x doesn't quite make it nice and simple

OpenStudy (anonymous):

-x(2x+1) look to see if 2x+1 is a factor of the bottom

OpenStudy (anonymous):

i dont think it is.. right?

OpenStudy (anonymous):

you can check if you'd like

OpenStudy (anonymous):

ok. checking.

OpenStudy (anonymous):

nope... it's not

OpenStudy (anonymous):

nope...

OpenStudy (anonymous):

-1 is a root

OpenStudy (anonymous):

Other than partial fraction. is there any shorcut to do this integration ?

OpenStudy (anonymous):

i mean any other method. don't care if it's hard.

OpenStudy (anonymous):

not that i can think of right now

OpenStudy (anonymous):

is it a definite or indefinite integral?

OpenStudy (anonymous):

indefinite.

OpenStudy (anonymous):

if it is finite, i can just use calculator ;P

OpenStudy (anonymous):

hmmm...

OpenStudy (anonymous):

see what you get after factoring out the root at -1. partial fraction decomp may be all you have. it's not that bad to use. have you learned it already?

OpenStudy (anonymous):

i think doing factoring with that imaginary root will make this problem more complex.

OpenStudy (anonymous):

i think, it's better i skip this question.

OpenStudy (anonymous):

no... you leave that as a quadratic if you get a complex root

OpenStudy (unklerhaukus):

but, partial fractions is fun

OpenStudy (anonymous):

it really is kind of cool, but takes a little bit of work

OpenStudy (anonymous):

it's ok. i think i will leave this question.. by the way 2 of you are helping me.. but i have to give pgpilot326 as he's trying hard to help me.

OpenStudy (anonymous):

thx for that... i'll see what i come up with on the decomp

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\frac{ -2x ^{2}-x }{ 2x ^{3}+x ^{2}+1 }=\frac{ -\frac{ 1 }{ 4 } }{ x+1 }+\frac{ -\frac{ 3 }{ 2 } x+\frac{ 1 }{ 4 }}{ 2x^2 - x+1 }\]